# Few-shot Class-incremental Learning: A Survey

### v2

@@ 1 Introduction
Figure 1: IL Publications from 2016 to 2023. It is observed that CIL research has become predominant in the field of IL over time, due to its practical value. Concurrently, FSCIL shows a steady rise, mirroring growing requirement of CIL with limited data.
@@
Although there has been some progress in the field of FSCIL and some representative works [15, 13, 14, 5, 16] have emerged, it is yet in its development stage. Current methods still have a gap to meet the practical applications. Therefore, it is imperative to systematically review the latest developments in this field, identify the core challenges and open questions that hinder its development, and determine the promising future direction. Nevertheless, most of the research on FSCIL is still quite dispersed, and this field needs a systematic and comprehensive survey. It has inspired our survey, which aims to fill the gap. Since it is an ML problem proposed in the field of computer vision in recent years and most of the research work is based on the deep learning algorithm, the scope discussed in our paper is mainly the deep FSCIL algorithm in the field of computer vision, which includes primarily classification and object detection tasks.
Despite existing surveys on FSL [17, 18, 19, 20] and IL [6, 21, 5, 22, 7, 23, 11, 24, 25], there is a clear lack of systematic and comprehensive surveys specifically on FSCIL. In this regard, we summarize existing surveys, list their characteristics in Tab. I, systematically describe the uniqueness of our paper to highlight its unique contributions. To address the shortcomings in FSCIL research, we systematically summarize the field from various aspects, including definition, challenges, general schemes, related problems, datasets, metrics, methods, performance comparisons, and future directions. Our contributions are:
@@
This paper provides a clear introduction of FSL, including definition, challenges, and main methods from data, model, and algorithm perspectives. However, it lacks a summary of datasets and performance.
2020 arXiv
Learning from Very Few Samples: A Survey [18]
@@ A Survey of Self-Supervised and Few-Shot Object Detection [20]
This paper reviews continual learning, highlights the need for stable real-world algorithms, summarizes existing benchmarks and metrics, and proposes a framework for evaluating approaches across both robotics and non-robotics fields. But it lacks a summary of performance.
@@ A Comprehensive Study of Class Incremental
This paper focuses on class incremental learning, providing the taxonomy of regularization approaches, dynamic architectures, and complementary learning systems and memory replay. However, it lacks a summary of challenges and relevant problems.
@@ A Continual Learning Survey: Defying Forgetting in Classification Tasks [5]
This paper summarizes task incremental learning and compares 11 advanced methods. It also analyzes the impact of model parameters, task order, and resource requirements. However, it lacks a summary of challenges and relevant problems.
@@
This study reviews online continual learning in image classification and compare advanced methods. However, it lacks a summary of challenges and relevant problems.
2023 arXiv
@@ A Comprehensive Survey of Continual Learning:
2023 arXiv
Deep Class-Incremental Learning: A Survey [24]
The paper categories methods into data-centric, model-centric, and algorithm-centric, providing the evaluation of 16 methods. The paper also argues for a fair comparison to avoid bias. But it lacks a summary of challenges and relevant problems.
@@ A Survey on Few-Shot Class-Incremental
This survey focuses on FSCIL. However, the taxonomy and summary of relevant problems are not systematic. It also lacks a summary of challenges. Discussion about object detection and the future direction is not in-depth.
@@
Our paper focuses on FSCIL in classification and object detection tasks, clarifies the challenge and relationship with relevant problems, provides a systemical taxonomy from data-based, structure-based, and optimization-based approaches. We also summarize the performance of existing methods and point out potential directions.
@@
Figure 2: A chronological overview of some representative FSCIL methods. FSCIL was first carried out by TOPIC [13]. CEC [14] was widely used as a base for subsequent studies. SaKD [27] integrated semantic word vectors into FSCIL, offering a reference for applying language-image models in the future. LCwoF [28] and ERDFR [29] proposed distinct data replay strategies. F2M [30] introduced a novel approach by constraining optimization within flat local minima. FSLL+SS [31] introduced the semi-supervised features to FSCIL for the first time. MgSvF [32] analyzed and utilized different frequency components to balance the old and new knowledge. FACT [33] introduced a fresh perspective by advocating forward compatibility in FSCIL. C-FSCIL [34] pre-defined classifiers to guide model optimization. DSN [35] offered a novel dynamic structure for FSCIL. LIMIT [16] proposed a representative meta-learning paradigm for FSCIL. CLOM [36] pointed out the issue of class-level overfitting made by metric learning in FSCIL.
@@ 2.1 Problem Definition
Figure 3: The general settings of different IL tasks. Specifically, (a) shows the setting of FSCIL, (b) is the CIL, (c) represents the setting of Task-incremental Learning (TIL), and (d) illustrates the Domain-incremental Learning (DIL). FSCIL can be viewed as a subdomain of CIL, where the base session usually has sufficient training data, and the incremental sessions are formed in the $N-$way $K-$shot format. TIL differs from CIL because the session identity is known during model training and testing. In contrast, DIL maintains the same classification tasks, but the data across different sessions comes from different domains. Note that “session” may also be called “task” in other literature.
@@
Model: During the training session $b$, only the dataset $D^{b}_{train}$ is accessible, as original complete training datasets from previous sessions are unavailable. The FSCIL model must learn new classes from $D^{b}_{train}$ while maintaining performance on old classes, i.e., minimizing the expected risk $\mathcal{R}\left(f,b\right)$ on all the seen classes [16, 5]. This process is formulated as follows:
@@ 2.2 Core Challenges
FSCIL faces significant challenges, notably unreliable empirical risk minimization and the stability-plasticity dilemma. In FSCIL sessions, limited supervised data mean empirical risk fails to accurately represent expected risk, decreasing model generalization and increasing overfitting risks. Moreover, as new classes are continually added, old knowledge can be easily forgotten and overwritten by new knowledge. This leads to catastrophic forgetting. Otherwise, intransigence may occur. Therefore, balancing model stability and plasticity is another core challenge. This section provides the details of these challenges.
@@ 2.2.1 Unreliable Empirical Risk Minimization
$\mathbb{E}\left[\mathcal{R}\left(f_{e}\right)-\mathcal{R}\left(f_{o}\right)\right]=\underbrace{\mathbb{E}\left[\mathcal{R}\left(f_{b}\right)-\mathcal{R}\left(f_{o}\right)\right]}_{\mathcal{E}_{app}}+\underbrace{\mathbb{E}\left[\mathcal{R}\left(f_{e}\right)-\mathcal{R}\left(f_{b}\right)\right]}_{\mathcal{E}_{est}},$ (7)
Here, the expectation concerns the random choice of $D_{train}$. The approximation error $\mathcal{E}{app}$ measures how closely functions in $\mathcal{F}$ can approximate the optimal hypothesis $f_{o}$, and the estimation error $\mathcal{E}{est}$ measures the effect of minimizing the empirical risk $\mathcal{R}I(f)$ instead of the expected risk $\mathcal{R}(f)$ in $\mathcal{F}$. Overall, the hypothesis space $\mathcal{F}$ and the number of examples in $D_{train}$ affect the total error [17].
@@ 2.2.2 Stability-plasticity Dilemma
In FSCIL, a central challenge is the stability-plasticity dilemma, which involves balancing the model’s consistent performance on learned classes (stability) and its adaptability to new classes with limited samples (plasticity). Traditional deep learning models are typically static and can only handle previously learned classes. FSCIL demands continual learning of new classes with only a few available labeled training samples and without access to the original complete training data of old classes. It requires the model to maintain the stability of previously learned knowledge and plasticity in learning new knowledge. However, because the optimization objectives for old and new classes are different and conventional deep learning algorithms indiscriminately optimize the existing model parameters when learning new classes, the decision boundary usually shifts towards the new classes. Therefore, the decision boundary may show poor performance or even complete failure in previous classes. The phenomenon is known as catastrophic forgetting. Similarly, an excessive focus on maintaining the stability of the model’s old knowledge may limit its ability to learn new tasks, known as intransigence. Thus, achieving an equilibrium between stability and plasticity is pivotal in FSCIL.
@@ 2.3 General Schemes
FSCIL has two main frameworks, which are presented in Fig. 6. The first one comprises a feature extractor and a softmax classifier, while the second involves a feature embedding network and the nearest class mean classifier [32, 42, 43]. The entire network is trainable throughout the IL process in the first one. To counteract catastrophic forgetting, some studies [44, 45, 46] employ KD techniques to train models, thereby enabling them to maintain competent classification capabilities on previous classes while accommodating new ones. The second framework focuses on training a feature embedding network, which maps samples into a feature space where the distances between them represent semantic differences. This is followed by classification using the nearest class mean classifier. For instance, some studies [47] employ metric loss for the training of the embedding network, enabling it to learn more discriminative features and better adapt to incremental classes.
Figure 6: The general schemes of FSCIL. (a) “Feature Extractor + Softmax Classifier” has a trainable backbone with KD-based methods being the most typical method. (b) is “Feature Embedding + Nearest Mean Classifier” with a fixed backbone after base training. The most typical approach is feature embedding methods with cosine similarity classifier.
@@ 2.4.1 Incremental Learning
This section reviews the relationship and distinctions between FSCIL and other IL scenarios, specifically CIL, TIL, and DIL, as outlined by Van de Ven et al. [11]. Notably, the term ”session” in FSCIL is adopted here instead of the ”task” terminology used in other IL literature.
@@ 2.5 Taxonomy
Few-shot Class-incremental Learning Few-shot Class-incremental Classification (Section 4)Data-based Approaches (Section 4.1)Data Replay-based Methods (Section 4.1.1)Raw Replay-based MethodsGenerative Replay-based Methods Pseudo Scenarios-based Methods (Section 4.1.2)Pseudo Class-based Methods Pseudo Session-based Methods Structure-based Approaches (Section 4.2)Dynamic Structure-based Methods (Section 4.2.1)Graph-based Methods Other Methods Attention-based Methods (Section 4.2.2)Optimization-based Approaches (Section 4.3)Representation Learning-based Methods (Section 4.3.1)Metric Learning-based Methods Feature Space-based Methods Feature Fusion-based Methods Knowledge Distillation Methods (Section 4.3.2)Knowledge Distillation-based Methods with Balanced Data Optimized Knowledge Distillation-based Methods Meta learning Methods (Section 4.3.3)Other Methods (Section 4.3.4)Few-shot Class-incremental Object Detection (Section 5)Anchor-free approaches (Section 5.2.1)CentreNet-based Methods FCOS-based Methods Anchor-based approaches (Section 5.2.2)Mask RCNN-based Methods DETR-based Method
@@ 3 Datasets and Evaluation
This section focuses on the role of datasets in FSCIL, crucial for benchmarking and comparing algorithms. It begins with a review of the primary public datasets for FSCIL, detailing their characteristics, statistics, and implementation specifics. Subsequently, it introduces evaluation metrics for various tasks, and concludes with an overview of limitations and future trends.
@@ 3.2.1 Evaluation Metrics for Classification
In FSCIL, the model must learn new classes while retaining knowledge of previous ones. After each incremental session, it undergoes testing on a joint dataset encompassing all classes encountered thus far, with classification accuracy as the standard evaluation metric. The model is required to balance maintaining old class recognition with learning new ones. Moreover, post-completion of all incremental sessions, the model’s overall performance is assessed using various metrics. Common metrics include Performance Dropping (PD) rate and Average Accuracy (AA). PD rate measures the accuracy difference between the final incremental and base sessions, where lower values denote better FSCIL performance. AA calculates the mean accuracy across the base and all incremental sessions, with higher values indicating superior FSCIL performance.
@@ 3.2.2 Evaluation Metrics for Object Detection
In FSCIOD tasks, two approaches incorporate new incremental data: batch and continuous IL. Batch IL entails learning all new classes at once, while continuous IL adds new classes progressively. The more common setting is batch IL, akin to a single-session FSCIL. The predominant performance metric here is Average Precision (AP), which assesses the detection model’s effectiveness. AP is calculated separately for base classes, new classes, and all classes, with higher AP values across all classes indicating better FSCIOD performance. Additionally, some studies use Average Recall (AR) and AP50 as complementary metrics for a more comprehensive evaluation.
@@ 4 Few-shot Class-incremental Classification
Figure 8: The representative classification methods in FSCIL: (a) LCwoF, storing raw samples for replay to calibrate performance; (b) ERDFR, which synthesizes old samples for replay via Generative Adversarial Network (GAN) model; (c) FACT, ensuring model forward compatibility by creating virtual classes during base training; (d) CEC, transferring contex between new and old prototypes using a GAT; and (e) DSN, facilitating FSCIL with a dynamic structure that allows for expansion and autonomous compression.

### v3

# Few-shot Class-incremental Learning for Classification and Object Detection: A Survey

@@ 1 Introduction
Figure 1: IL publications from 2016 to 2023. It is observed that CIL research has become predominant in the field of IL over time, due to its practical value. Concurrently, FSCIL shows a steady rise, mirroring the growing requirement of CIL with limited data.
@@
Although there has been some progress in the field of FSCIL and some representative works [15, 13, 14, 5, 16] have emerged, it is yet in its development stage. The key milestones from 2020 to the present are illustrated in Fig. 2. Current methods still have a gap to meet the practical applications. Therefore, it is imperative to systematically review the latest developments in this field, identify the core challenges and open questions that hinder its development, and determine the promising future direction. Nevertheless, most of the research on FSCIL is still quite dispersed, and this field needs a systematic and comprehensive survey. It has inspired our survey, which aims to fill the gap. Since it is an ML problem proposed in the field of computer vision in recent years and most of the research work is based on the deep learning algorithm, the scope discussed in our paper is mainly the deep FSCIL algorithm in the field of computer vision, which includes primarily classification and object detection tasks.
Figure 2: A chronological overview of some representative FSCIL methods. FSCIL was first carried out by TOPIC [13]. CEC [14] was widely used as a base for subsequent studies. SaKD [17] integrated semantic word vectors into FSCIL, offering a reference for applying language-image models in the future. LCwoF [18] and ERDFR [19] proposed distinct Data Replay (DR) strategies. F2M [20] introduced a novel approach by constraining optimization within flat local minima. FSLL+SS [21] introduced the semi-supervised features to FSCIL for the first time. MgSvF [22] analyzed and utilized different frequency components to balance the old and new knowledge. FACT [23] introduced a fresh perspective by advocating forward compatibility in FSCIL. C-FSCIL [24] pre-defined classifiers to guide model optimization. DSN [25] offered a novel dynamic structure for FSCIL. LIMIT [16] proposed a representative meta-learning paradigm for FSCIL. CLOM [26] pointed out the issue of class-level overfitting made by metric learning in FSCIL.
Despite existing surveys on FSL [27, 28, 29, 30] and IL [6, 31, 5, 32, 7, 33, 11, 34, 35], there is a clear lack of systematic and comprehensive surveys specifically on FSCIL. Existing FSL surveys primarily focus on tackling ML problems with limited data. While they provide systematic classifications of FSL methods, they do not touch upon the issue of FSCIL. Similarly, IL surveys mostly focus on ML problems in a continual learning scenario. For instance, De Lange et al. [7] conducted a systematic review of DR, regularization, and parameter isolation, along with comparative experiments, while Zhou et al. [34] summarized the CIL problem from perspectives like DR, Data Regularization, and Dynamic Networks. However, none of these works systematically review FSCIL. Although some surveys briefly mention FSCIL [32, 35], they only provide a short introduction to the concept and a few studies without offering a thorough or systematic analysis. Although Tian et al. [36] recently conducted a review of FSCIL, its introduction to FSCIL is not in-depth and comprehensive enough.
In this regard, we summarize existing surveys in Tab. I and systematically describe the uniqueness of our paper to highlight its unique contributions. To address the shortcomings in FSCIL research, we systematically summarize the field from various aspects, including definition, challenges, general schemes, related problems, datasets, metrics, methods, performance comparisons, and future directions. Our contributions are:
@@
This paper clearly introduces FSL, including its definition, challenges, and main methods from data, model, and algorithm perspectives. However, it lacks a summary of datasets and performance.
@@ A Survey of Self-Supervised and Few-Shot Object Detection [30]
This paper reviews continual learning, highlights the need for stable real-world algorithms, summarizes existing benchmarks and metrics, and proposes a framework for evaluating approaches across both robotics and non-robotics fields. However, it lacks a summary of performance.
@@ A Comprehensive Study of Class Incremental
This paper focuses on CIL, providing the taxonomy of regularization approaches, dynamic architectures, and complementary learning systems and memory replay. However, it lacks a summary of challenges and relevant problems.
@@ A Continual Learning Survey: Defying Forgetting in Classification Tasks [5]
This paper summarizes Task-incremental Learning (TIL) and compares 11 advanced methods. It also analyzes the impact of model parameters, task order, and resource requirements. However, it lacks a summary of challenges and relevant problems.
@@
This study reviews online continual learning in image classification and compares advanced methods. However, it lacks a summary of challenges and relevant problems.
2024 TPAMI
@@ A Comprehensive Survey of Continual Learning:
2024 TPAMI
Class-Incremental Learning: A Survey [34]
This paper categorizes methods into data-centric, model-centric, and algorithm-centric, evaluating 16 methods and advocating for fair comparisons. However, it lacks a summary of challenges and relevant problems.
@@ A Survey on Few-Shot Class-Incremental
This survey focuses on FSCIL. However, the taxonomy and summary of relevant problems are not systematic. It also lacks a summary of the challenges. Discussion about object detection and the future direction is not in-depth.
@@
Our paper focuses on FSCIL in classification and object detection tasks, clarifies the challenge and relationship with relevant problems, and provides a systemical taxonomy from data-based, structure-based, and optimization-based approaches. We also summarize the performance of existing methods and point out potential directions.
@@ 2.1 Problem Definition
Figure 3: The general settings of different IL tasks. Specifically, (a) shows the setting of FSCIL, (b) is the CIL, (c) represents the setting of TIL, and (d) illustrates the Domain-incremental Learning (DIL). FSCIL can be viewed as a subdomain of CIL, where the base session usually has sufficient training data, and the incremental sessions are formed in the $N-$way $K-$shot format. TIL differs from CIL because the session identity is known during model training and testing. In contrast, DIL maintains the same classification tasks, but the data across different sessions comes from different domains. Note that “session” may also be called “task” in other literature.
@@
Model: During the training session $b$, the dataset $D^{b}_{train}$ is accessible, and the original complete training datasets from previous sessions are unavailable (note: some methods like DR or KD may store a few old samples to revisit old knowledge). The FSCIL model must learn new classes from $D^{b}_{train}$ while maintaining performance on old classes, i.e., minimizing the expected risk $\mathcal{R}\left(f,b\right)$ on all the seen classes [16, 5]. This process is formulated as follows:
@@ 2.2 Core Challenges
FSCIL faces significant challenges, notably the unreliable empirical risk minimization and the stability-plasticity dilemma. In FSCIL sessions, limited supervised data mean empirical risk fails to accurately represent expected risk, decreasing model generalization and increasing overfitting risks. Moreover, as new classes are continually added, old knowledge can be easily forgotten and overwritten by new knowledge. This leads to catastrophic forgetting. Otherwise, intransigence may occur. Therefore, balancing model stability and plasticity is another core challenge. This section provides the details of these challenges.
@@ 2.2.1 Unreliable Empirical Risk Minimization
$\mathbb{E}\left[\mathcal{R}\left(f_{e}\right)-\mathcal{R}\left(f_{o}\right)\right]=\underbrace{\mathbb{E}\left[\mathcal{R}\left(f_{b}\right)-\mathcal{R}\left(f_{o}\right)\right]}_{\mathcal{E}_{app}}+\underbrace{\mathbb{E}\left[\mathcal{R}\left(f_{e}\right)-\mathcal{R}\left(f_{b}\right)\right]}_{\mathcal{E}_{est}}.$ (7)
Here, the expectation concerns the random choice of $D_{train}$. The approximation error $\mathcal{E}_{app}$ measures how closely functions in $\mathcal{F}$ can approximate the optimal hypothesis $f_{o}$, and the estimation error $\mathcal{E}_{est}$ measures the effect of minimizing the empirical risk $\mathcal{R}_{I}(f)$ instead of the expected risk $\mathcal{R}(f)$ in $\mathcal{F}$. Overall, the hypothesis space $\mathcal{F}$ and the number of examples in $D_{train}$ affect the total error [27].
@@ 2.2.2 Stability-plasticity Dilemma
In FSCIL, a central challenge is the stability-plasticity dilemma, which involves balancing the model’s consistent performance on learned classes (stability) and its adaptability to new classes with limited samples (plasticity). Traditional deep learning models are typically static and can only handle previously learned classes. FSCIL demands continual learning of new classes with only a few available labeled training samples and without access to the original complete training data of old classes. It requires the model to maintain the stability of previously learned knowledge and plasticity in learning new knowledge. Due to different optimization goals for old and new classes, the decision boundary often shifts toward new classes, leading to catastrophic forgetting. Conversely, focusing too much on old knowledge stability may limit the ability to learn new tasks, a phenomenon known as intransigence. Therefore, balancing stability and plasticity is crucial in FSCIL.
@@
Figure 6: The general schemes of FSCIL. (a) “Feature Extractor + Softmax Classifier” has a trainable backbone with Knowledge Distillation (KD)-based methods being the most typical method. (b) is “Feature Embedding + Nearest Mean Classifier” with a fixed backbone after base training.
@@ 2.3 General Schemes
FSCIL has two main frameworks, as shown in Fig. 6. The first uses a feature extractor with a softmax classifier, while the second involves a feature embedding network and the nearest class mean classifier [22, 42, 43]. The entire network is trainable throughout the IL process in the first one. To mitigate catastrophic forgetting, some studies [44, 45, 46] use KD to maintain competent classification capabilities on previous classes while learning new ones. The second framework focuses on training a feature embedding network to map samples into a space where distances represent semantic differences, followed by classification using the nearest class mean classifier. For instance, some studies [47] employ metric loss for the training of the embedding network, enabling it to learn more discriminative features and better adapt to incremental classes.
@@ 2.4.1 Incremental Learning
This section reviews the relationship and distinctions between FSCIL and other IL scenarios, specifically CIL, TIL, and DIL, as outlined by Van de Ven et al. [11].
@@ 2.5 Taxonomy
Few-shot Class-incremental Learning Few-shot Class-incremental Classification (Section 4)Data-based Approaches (Section 4.1)Data Replay-based Methods (Section 4.1.1)Raw Replay-based MethodsGenerative Replay-based Methods Pseudo Scenarios-based Methods (Section 4.1.2)Pseudo Class-based Methods Pseudo Session-based Methods Structure-based Approaches (Section 4.2)Dynamic Structure-based Methods (Section 4.2.1)Graph-based Methods Other Methods Attention-based Methods (Section 4.2.2)Optimization-based Approaches (Section 4.3)Representation Learning-based Methods (Section 4.3.1)Metric Learning-based Methods Feature Space-based Methods Feature Fusion-based Methods Knowledge Distillation Methods (Section 4.3.2)Knowledge Distillation-based Methods with Balanced Data Optimized Knowledge Distillation-based Methods Meta learning Methods (Section 4.3.3)Other Methods (Section 4.3.4)Few-shot Class-incremental Object Detection (Section 5)Anchor-free approaches (Section 5.2.1)CentreNet-based Methods FCOS-based Methods DETR-based Method Anchor-based approaches (Section 5.2.2)Mask RCNN-based Methods
@@ 3.2.1 Evaluation Metrics for Classification
In FSCIL, the model needs to learn new classes while retaining previous knowledge. After each session, it is tested on all encountered classes, using accuracy as the primary metric. Additionally, after completing all incremental sessions, the overall performance is evaluated using Average Accuracy (AA) and Performance Dropping (PD) rate. The AA calculates the mean accuracy across the base and all incremental sessions, with higher values indicating superior performance. PD measures the accuracy drop between the base and final incremental sessions, where lower values represent better FSCIL performance. Definitions are shown in Tab. III.
TABLE III: Evaluation metric definitions for classification and object detection tasks. $A_{i}$ represents the accuracy obtained in session $i$, and $B$ is the number of incremental sessions in classification. $P_{i}$ and $R_{i}$ represent the precision and recall obtained in class $i$, and $K$ is the number of counted classes.
Task Metric
Classification $AA=\frac{1}{B+1}\sum_{i=0}^{B}A_{i}$ $PD=A_{B}-A_{0}$
Object Detection $mAP=\frac{1}{K}\sum_{i=1}^{K}P_{i}$ $mAR=\frac{1}{K}\sum_{i=1}^{K}R_{i}$
@@ 3.2.2 Evaluation Metrics for Object Detection
In FSCIOD tasks, two approaches incorporate new incremental data: batch and continuous IL. Batch IL entails learning all new classes at once, while continuous IL adds new classes progressively. Batch IL, similar to single-session FSCIL, is more common. The predominant performance metric is mean Average Precision (mAP), which is the mean of the AP values calculated for all counted classes. mAP is calculated separately for base classes, new classes, and all classes, with higher mAP values across all classes indicating better FSCIOD performance. Additionally, some studies use a similar way to calculate mean Average Recall (mAR) and mAP50 as complementary metrics for a more comprehensive evaluation. Definitions are shown in Tab. III.
@@ 4 Few-shot Class-incremental Classification
Figure 8: The representative classification methods in FSCIL, with the core designs highlighted in red. (a) LCwoF stores some old samples for raw replay, jointly calibrating performance with the new classes during new sessions; (b) ERDFR trains a generator using the previous main model and synthesizes virtual old samples for DR during the learning of new classes, aiding in the retention of old knowledge; (c) FACT creates virtual incremental classes from base classes to simulate future scenarios, enabling the model to develop forward compatibility during the base session; (d) CEC constructs a pseudo-incremental session from the base session to train a GAT, which is later used to adjust the relationship between new and old class prototypes during real incremental sessions; (e) DSN designs a dynamic structure alongside the backbone that allows for expansion and autonomous compression, facilitating the learning of new classes while retaining old knowledge.